OpenStudy (macaylam):

Which ordered pair (p, r) is a solution to the given system of linear equations? 5p-3r=1 8p+6r=4 A.(–1, –2) B. (–2, –1) C.(2/9,1/3) D.(1/3,2/9)

OpenStudy (gunsodo):

Easiest way : Put the value from the choices into two equations

OpenStudy (gunsodo):

Best way : Eliminate a variable from the system of equation, by multiplying one with a constant, to make the coefficient of a variable equal. Then eliminate

OpenStudy (macaylam):

is it a?

OpenStudy (wanell):

5p - 3r = 1 /*2 8p + 6r = 4 __________ 10p - 6r = 2 8p + 6r = 4 __________ 10p - 6r + 8p + 6r = 2 + 4 18p = 6 p = 6/18 p = 1/3 5p - 3r = 1 5 * 1/3 - 3r = 1 5/3 - 3r = 1 5/3 - 1 = 3r 5/3 - 3/3 = 3r 2/3 = 3r r = 2/9 (p, r) = (1/3, 2/9) The correct result would be (1/3, 2/9).

OpenStudy (macaylam):

thank you